# Around the Merino–Welsh conjecture: improving Jackson’s inequality

### v1

@@ 4. The main lemma
$G(d,x,s,\gamma_{x,s}(1))\geq 1.$
In order to check the first inequality for large $d$ it is enough if for some $d_{0}$ we have
$G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geq 1,$
then we get that
$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{\min(2,d-1)}\geq 1$
for $d\geq d_{0}$. So we only need to check the first inequality for $d=2,\dots,d_{0}$ and the inequality $G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geq 1$. These inequalities are satisfied for $x=2.36$, $s=0.78$ and $d_{0}=44$. Again see the jupyter notebook [7] or alternatively, Table 3.
Idea 4. We can gain a very small further improvement as follows. For $d\geq 2$ we can use that
$\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)$
$\geq\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\gamma_{x,s}(1)^{\min(2,d_{v}-1)}\gamma_{x,s}(2)^{d_{v}-\min(2,d_{v}-1)}\right),$
@@
$G_{2}(d,x,s,\gamma_{1},\gamma_{2}):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma_{1}^{\min(2,d-1)}\gamma_{2}^{d-\min(2,d-1)}\right).$
Then it is enough to check that
$G_{2}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\geq 1$
for $d\geq 2$ and
$G_{2}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))G(d,x,s,\gamma_{x,s}(1))^{\min(2,d-1)}\geq 1$
for all $d\geq 2$. This leads to the tiny improvement $x=2.355$ ($s=0.78$ and $d=100$). See [7] or Table 4 in the Appendix for further details.
@@ 6. Matroids with prescribed circuit lengths
Let $k\geq 4$ be a real number. Let $H$ be a bipartite graph such that all degrees of $H$ lie in the interval $[k+1,k^{4}-2k^{2}-1]$. Then
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
@@
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\prod_{v\in V}G(d_{v},2,s,s).$
@@
We will show that if $s=1-\frac{1}{k^{2}}$, then for all $d\in[k+1,k^{4}-2k^{2}-1]$ we have $G(d,2,s,s)\geq 1$. First observe that $G(d,2,s,s)$ is a concave function of $s$ on the interval $(0,1)$ as its second derivative is
$-\frac{(2-s)(d^{2}+d)s^{d}+2ds}{(d+1)s},$
which is clearly negative. This means that $I_{d}:=\{s\ |\ G(d,2,s,s)\geq 1\}$ is an interval for each $d$. We will show that $1-\frac{1}{k^{2}}\in I_{k+1}\cap I_{k^{4}-2k^{2}-1}$. By first choosing $k=d-1$ we get that $1-\frac{1}{(d-1)^{2}}\in I_{d}$. If we choose $k^{2}=1+\sqrt{d+2}$, then $k^{4}-2k^{2}-1=d$ and so $1-\frac{1}{1+\sqrt{d+2}}\in I_{d}$. Hence $\left[1-\frac{1}{1+\sqrt{d+2}},1-\frac{1}{(d-1)^{2}}\right]\subseteq I_{d}$. This means that if $k+1\leq d\leq k^{4}-2k^{2}-1$, then $1-\frac{1}{k^{2}}\in I_{d}$.
First we show that $1-\frac{1}{k^{2}}\in I_{k+1}$, that is, $G(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}})\geq 1$. We will use that
$(1-t)^{r}\leq 1-\binom{r}{1}t+\binom{r}{2}t^{2}$
for $0\leq t\leq 1$ by Bonferroni’s inequality applied to independent events with probability $t$. Hence
@@
$\geq\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{d+1}{k^{2}}+\frac{(d+1)d}{2k^{2}}\right)\right)$
@@
$G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $\geq\left(\frac{k+3}{k+2}+\frac{k+1}{k+2}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{k+2}\left(1-\frac{k+2}{k^{2}}+\frac{(k+2)(k+1)}{2k^{2}}\right)\right)$
$=1+\frac{1}{2}\frac{3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2}{k^{8}+4k^{7}+4k^{6}}$
For $k\geq 4$ we have $03k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2>$ 0 showing that $G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)>1$.
@@
$\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}<e^{-\frac{1}{k^{2}}(k^{4}-2k^{2})}=e^{-k^{2}-2}.$
Let us introduce $y=k^{2}-2$. Then $k^{6}-k^{4}-k^{2}-1=y^{3}+5y^{2}+7y+1$, so we need that
$\frac{y}{y^{3}+5y^{2}+7y+1}<e^{-y}$
@@
Since $k\geq 3$ we have $y=k^{2}-2\geq 7$, but $e^{y}>y^{2}+5y+7+\frac{1}{y}$ already true for $y\geq 4$.
Remark 6.2.
There was no attempt to optimize Lemma 6.1, not even concerning the minimum degree. It turns out that if $s=0.9226$, then
$G(d,2,s,\gamma_{2,s}(3))>1$
for $3\leq d\leq 141$. So if all degrees of a bipartite graph $H$ lies in the interval $[3,141]$, then
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
Similarly, for $s=0.9622$ we have
$G(d,2,s,\gamma_{2,s}(4))>1$
for $4\leq d\leq 646$. So if all degrees of a bipartite graph $H$ lies in the interval $[4,646]$, then
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
This shows that in Theorem 1.4 we could have written $\ell\geq 4$ and probably a faster growing function than $(\ell-2)^{4}$.
@@
Let $A$ be a basis of the matroid $M$ on ground set $E$. Let $B=E\setminus A$. Let us consider the local basis exchange graph $H[A]$ on $A\cup B$. Observe that $v\in B$ are adjacent with exactly those elements that are in his fundamental circuit. So $\ell\leq d_{v}+1\leq(\ell-2)^{4}$. Similarly, if $u\in A$, then it is adjacent with those elements of $B$ that are in its fundamental circuits with respect to $B$ in $M^{*}$. So $\ell\leq d_{u}+1\leq(\ell-2)^{4}$.
By choosing $k=\ell-1$ we get that $k^{4}-2k^{2}-1=(\ell-1)^{4}-2(\ell-1)^{2}-1\geq(\ell-2)^{4}$ if $\ell\geq 3$. So by Lemma 1.4 we have
$\widetilde{T}_{H[A]}(2,0)\widetilde{T}_{H[A]}(0,2)\geq\widetilde{T}_{H[A]}(1,1)^{2}.$
Since this is true for all basis $A$ of the matroid $M$, we get by the Transfer lemma (Lemma 2.6 and Remark 2.7) that
$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
@@ 7. Concluding remarks
We end this paper with some conjectures.
@@
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geq\widetilde{T}_{H}(1,1)^{2}.$
If this conjecture is true, then the following conjecture would be an immediate corollary.
@@
If $M$ is a simple co-simple matroid, then
$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
Acknowledgment. The author is very grateful to Csongor Beke, Gergely Kál Csáji and Sára Pituk for discussions on the topic of this paper.

### v2

@@ 4. The main lemma
$G(d,x,s,\gamma_{x,s}(1))\geqslant G(\infty,x,s)\geqslant 1.$
With the choice $x=2.36,s=0.78$ and $d_{0}=44$ all these inequalities are satisfied. Then we can check the remaining inequalities for $d=3,\dots,43$ with a computer separately. Again see the jupyter notebook [7] or alternatively, Table 3.
Idea 4. We can gain a very small further improvement as follows. We start our series of inequalities just as in the previous step:
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}\mathbb{G}\left(d_{u},x,s,\{\gamma_{x,s}(d_{u^{\prime}})\}_{u^{\prime}\in N(u)}\right)\right)$
$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
Previously, we bounded $\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$ with $G(d_{v},x,s,\gamma_{x,s}(1))$, but now we observe that if $d_{v}>2$, then $v$ can have at most two neighbors of degree $1$, while if $d_{v}=2$ then $v$ can have at most one vertex of degree $1$, all other neighbors have degree at least $2$. Thus, for $d_{v}\geq 2$ we can use that
$\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
$\geqslant\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\gamma_{x,s}(1)^{\min(2,d_{v}-1)}\gamma_{x,s}(2)^{d_{v}-\min(2,d_{v}-1)}\right),$
@@
$G_{k}(d,x,s,\gamma_{1},\gamma_{2}):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma_{1}^{k}\gamma_{2}^{d-k}\right).$
Then it is enough to check that for all $d\geqslant 2$ we have
$\min_{0\leqslant k\leqslant\min(2,d-1)}G_{k}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\cdot G(1,x,s,\gamma_{x,s}(d))^{k}\geqslant 1.$
We show that this is indeed true if $x=2.355$ and $s=0.78$. First, we check it for $d\geqslant d_{0}=100$. By Lemma 4.6 the sequence $G(d,x,s,\gamma)$ is monotone decreasing for $d\geqslant d_{0}$ if $\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}\geqslant\gamma$. Thus for $0\leqslant k\leqslant 2$ we have
$G_{k}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\cdot G(1,x,s,\gamma_{x,s}(d))^{k}$ $\geqslant G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d_{0}))^{k}$
$\geqslant G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{k}$
$>1,$
where the last inequality can be checked directly. For $d<d_{0}$ we again check the values by a computer. See [7] or Table 4 in the Appendix for further details.
Remark 5.1.
It turns out that the value $2.355$ can be further improved to $2.35$ by observing that one can disregard one particular inequality:
$G_{2}(3,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))G(1,x,s,\gamma_{x,s}(3))^{2}\geqslant 1.$
This inequality comes from the situation where a vertex $v$ of degree $3$ is adjacent to two vertices of degree $1$. But this means that $H$ can be obtained by gluing $H_{1}=S_{4}$ to some $H_{2}$ at one of its leaves. As we discussed at Idea 3 this would mean that $H$ is not a minimal counter example. Choosing $x=2.35$, $s=0.79$ and $d_{0}=1000$ the remaining parts of Idea 4 work verbatimely.
@@ 6. Matroids with prescribed circuit lengths
Let $k\geqslant 4$ be a real number. Let $H$ be a bipartite graph such that all degrees of $H$ lie in the interval $k+1,k^{4}-2k^{2}-1]$. Then
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\widetilde{T}_{H}(1,1)^{2}.$
@@
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\prod_{v\in V}G(d_{v},2,s,s).$
@@
We will show that if $s=1-\frac{1}{k^{2}}$, then for all $d\in[k+1,k^{4}-2k^{2}-1]$ we have $G(d,2,s,s)\geqslant 1$. First, observe that $G(d,2,s,s)$ is a concave function of $s$ on the interval $(0,1)$ as its second derivative is
$2\frac{d}{d+1}(1-s^{d})+\left(\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)\right)(-ds^{d-1})$
which is clearly negative as both terms are negative. (Here we used the derivation rule $(h_{1}\cdot h_{2})^{\prime\prime}=h_{1}^{\prime\prime}\cdot h_{2}+2h_{1}^{\prime}\cdot h_{2}^{\prime}+h_{1}\cdot h_{2}^{\prime\prime}$ for $h_{1}=\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)$ and $h_{2}=s-\frac{s^{d+1}}{d+1}$ with $h_{1}^{\prime}=-\frac{d}{d+1}$, $h_{1}^{\prime\prime}=0$, $h_{2}^{\prime}=1-s^{d}$, $h_{2}^{\prime\prime}=-ds^{d-1}$.)
This means that
$I_{d}:=\{s\ |\ G(d,2,s,s)\geqslant 1\}$
is an interval for each $d$. We will show that $1-\frac{1}{k^{2}}\in I_{k+1}\cap I_{k^{4}-2k^{2}-1}$. By first choosing k=d1k=d-1 we get that $1-\frac{1}{(d-1)^{2}}\in I_{d}$. If we choose $k^{2}=1+\sqrt{d+2}$, then $k^{4}-2k^{2}-1=d$ and so $1-\frac{1}{1+\sqrt{d+2}}\in I_{d}$. Hence $\left[1-\frac{1}{1+\sqrt{d+2}},1-\frac{1}{(d-1)^{2}}\right]\subseteq I_{d}$. This means that if $k+1\leqslant d\leqslant k^{4}-2k^{2}-1$, then $1-\frac{1}{k^{2}}\in I_{d}$.
First we show that $1-\frac{1}{k^{2}}\in I_{k+1}$, that is, $G(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}})\geqslant 1$. We will use that
$(1-t)^{r}\leqslant 1-\binom{r}{1}t+\binom{r}{2}t^{2}$
for $0\leqslant t\leqslant 1$. This can be either proved by Bonferroni’s inequality applied to independent events with probability $t$, or by induction: the base case $r=1$ being trivial we have
$(1-t)^{r}$ $=(1-t)^{r-1}(1-t)$
$\leqslant\left(1-\binom{r-1}{1}t+\binom{r-1}{2}t^{2}\right)(1-t)$
$=1-\binom{r}{1}t+\binom{r}{2}t^{2}-\binom{r-1}{2}t^{2}$
$\leqslant 1-\binom{r}{1}t+\binom{r}{2}t^{2}.$
Hence
@@
$\geqslant\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{d+1}{k^{2}}+\frac{(d+1)d}{2k^{4}}\right)\right).$
@@
$G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $\geqslant\left(\frac{k+3}{k+2}+\frac{k+1}{k+2}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{k+2}\left(1-\frac{k+2}{k^{2}}+\frac{(k+2)(k+1)}{2k^{4}}\right)\right)$
$=1+\frac{1}{2}\frac{3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2}{k^{8}+4k^{7}+4k^{6}}.$
For $k\geqslant 4$ we have $3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2>0$ showing that $G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)>1$.
@@
$\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}<e^{-\frac{1}{k^{2}}(k^{4}-2k^{2})}=e^{-k^{2}+2}.$
Let us introduce $y=k^{2}-2$. Then $k^{6}-k^{4}-k^{2}-1=y^{3}+5y^{2}+7y+1$, so it is enough to prove that
$\frac{y}{y^{3}+5y^{2}+7y+1}>e^{-y}$
@@
Since $k\geqslant 3$ we have $y=k^{2}-2\geqslant 7$, but $e^{y}>y^{2}+5y+7+\frac{1}{y}$ is already true for $y\geqslant 4$.
@@
Let $A$ be a basis of the matroid $M$ on ground set $E$. Let $B=E\setminus A$. Let us consider the local basis exchange graph $H[A]$ on $A\cup B$. Observe that $v\in B$ are adjacent with exactly those elements that are in its fundamental circuit. So $\ell\leqslant d_{v}+1\leqslant(\ell-2)^{2}(\ell^{2}-4\ell+2)$. Similarly, if $u\in A$, then it is adjacent with those elements of $B$ that are in its fundamental circuit with respect to $B$ in $M^{*}$. So $\ell\leqslant d_{u}+1\leqslant(\ell-2)^{2}(\ell^{2}-4\ell+2)$.
We treat the cases $\ell=4$ and $5$ separately. If $\ell=4$, then all degree $d_{v}$ satisfy that $3\leqslant d_{v}\leqslant 7$. Let us choose $s=0.9226$. Then
$G(d,2,s,\gamma_{2,s}(3))>1$
for $3\leqslant d\leqslant 141$. By Corollary 4.5 we have
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(3))>1=\widetilde{T}_{H}(1,1)^{2}.$
Together with the transfer lemma this implies the claim. In case of $\ell=5$ all degree $d_{v}$ satisfy $4\leqslant d_{v}\leqslant 62$. With the choice $s=0.9622$ we have
$G(d,2,s,\gamma_{2,s}(4))>1$
for $4\leqslant d\leqslant 646$. Hence again by Corollary 4.5 we have
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(4))>1=\widetilde{T}_{H}(1,1)^{2}$
and the transfer lemma this implies the claim.
Now let us assume that $\ell\geqslant 6$. By choosing $k=\ell-2$ we get that $k\geq 4$ and for all vertex $v$ we have $k+1=\ell-1\leq d_{v}$ and
$k^{4}-2k^{2}-1=(\ell-2)^{4}-2(\ell-2)^{2}-1\geqslant(d_{v}+1)-1=d_{v}.$
This means that all degree $d_{v}$ satisfy $k+1\leqslant d_{v}\leqslant k^{4}-2k^{2}-1$. So by Lemma 6.1 we have
$\widetilde{T}_{H[A]}(2,0)\widetilde{T}_{H[A]}(0,2)\geqslant\widetilde{T}_{H[A]}(1,1)^{2}.$
Since this is true for all basis $A$ of the matroid $M$, we get by the Transfer lemma (Lemma 2.6) that
$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}.$
@@ 7. Concluding remarks
We end this paper with some conjectures. Let us start with an admittedly provocative conjecture.
Conjecture 7.1.
Let $x_{0}$ be the largest zero of the polynomial $x^{3}-9(x-1)$. Then for any matroid $M$ without loops and coloops we have
$T_{M}(x_{0},0)T_{M}(0,x_{0})\geq T_{M}(1,1)^{2}.$
We believe that even a disproof of this conjecture would lead to a better understanding of this problem than slightly improving the constant $2.355$ (or $2.35$).
Concerning the Merino-Welsh conjecture, the following variant might be true.
Conjecture 7.2.
If $M$ is a simple co-simple matroid, then
$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}.$
The following line of attack seems to be natural.
@@
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\widetilde{T}_{H}(1,1)^{2}.$
Clearly, Conjecture 7.3 implies Conjecture 7.2. At this point, we must admit that we lack examples. Computational results suggest the following conjecture. (Here we only phrase the conjecture when the number of vertices is divisible by $4$, but there are similar conjectures when $n=4k+i$ for $i\in\{1,2,3\}$.)
@@
Let $n=4k$, then the following bipartite graph minimizes $\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)$ among bipartite graphs on $n$ vertices without isolated vertices: add a leaf vertex to each vertex of a $K_{k,k}$.
Figure 3. The conjectured minimizer graph for $\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)$ on $24$ vertices.
If this conjecture is true, then the Merino-Welsh conjecture is true for matroids on $n$ elements when $n\leqslant 40$ and divisible by $4$, which shows why it is hard to find a counter example for the Merino-Welsh conjecture just by brute force computer search.
Acknowledgment. The author is very grateful to Csongor Beke, Ferenc Bencs, Gergely Kál Csáji and Sára Pituk for discussions on the topic of this paper. The author also thanks the very careful work of the two referees, they spotted many typos and gave very useful suggestions.